OpenStudy (anonymous):

Find three consecutive even integers whose sum is between 30 and 51.

Parth (parthkohli):

Let's say that the sum is 40. the three even integers would be x, x + 2 and x + 4. can you do it now?

OpenStudy (anonymous):

So it will be x+x+2+x+4=40

Parth (parthkohli):

Yes, can you solve now?

OpenStudy (anonymous):

3x+6=40

OpenStudy (anonymous):

3x=40-6

OpenStudy (anonymous):

X=12